# Byzantine Resilient Federated Multi-Task Representation Learning

### v2

@@ Abstract
In this paper, we propose BR-MTRL, a Byzantine-resilient multi-task representation learning framework that handles faulty or malicious agents. Our approach leverages representation learning through a shared neural network model, where all clients share fixed layers, except for a client-specific final layer. This structure captures shared features among clients while enabling individual adaptation, making it a promising approach for leveraging client data and computational power in heterogeneous federated settings to learn personalized models. To learn the model, we employ an alternating gradient descent strategy: each client optimizes its local model, updates its final layer, and sends estimates of the shared representation to a central server for aggregation. To defend against Byzantine agents, we employ two robust aggregation methods for client-server communication, Geometric Median and Krum. Our method enables personalized learning while maintaining resilience in distributed settings. We implemented the proposed algorithm in a federated testbed built using Amazon Web Services (AWS) platform and compared its performance with various benchmark algorithms and their variations. Through experiments using real-world datasets, including CIFAR-10 and FEMNIST, we demonstrated the effectiveness and robustness of our approach and its transferability to new unseen clients with limited data, even in the presence of Byzantine adversaries.
@@ I Introduction
Machine learning has traditionally relied on centralized settings, where a single model is trained on large datasets stored in one central location. This approach requires gathering data from all clients into a central data center, which raises significant data privacy concerns. Federated Learning (FL) addresses these challenges by using a distributed framework that preserves data privacy while overcoming the problem of isolated data sources [1]. FL enables multiple clients, such as mobile devices or institutions, to collaboratively train a global model without sharing their raw data. A central server is responsible for coordinating the clients by aggregating the models received from each client and broadcasting the updated model back to them in each round of communication. McMahan et al. introduced the Federated Averaging (FedAvg) algorithm, which facilitates the training of a single shared global model designed to achieve optimal performance on average across all participating clients [1]. The approach offers several advantages, such as simplicity in implementation and a streamlined aggregation process, making it scalable for large client populations. Additionally, the global model promotes uniformity, ensuring consistent performance for clients with similar data distributions. However, FL approaches encounter three primary bottlenecks: (i) Data heterogeneity: In real-world scenarios, client tasks often exhibit significantly different underlying data distributions. In such environments, a global model optimized for average loss may underperform across diverse clients, as it fails to account for these variations in data [2]. (ii) Generalization and adaptability: Federated models may struggle to generalize to unseen clients that were not part of the training process, limiting their effectiveness in dynamic, real-world applications [3]. (iii) Byzantine attacks: FL is particularly vulnerable to Byzantine attacks, as it relies on distributed client updates; a faulty or malicious client can potentially corrupt the global model, compromising performance and reliability [4].
@@
To solve data heterogeneity, many personalized FL algorithms have been developed. Yuyang Deng et al. proposed the Adaptive Personalized Federated Learning (APFL) algorithm to solve the problem of statistical heterogeneity in federated learning by applying a mix of local and global model optimization [5]. This method allows clients to train personalized models that better generalize on local distributions while still contributing to the learning of a robust global model. In addition, Alireza Fallah et al. proposed a method called Personalized Federated Learning via Model-Agnostic Meta-Learning (MAML) to solve the problem of personalized adaptation in federated learning by applying a meta-learning framework [6]. This method allows clients to quickly adapt a shared initial model to their local data distributions by performing a few steps of gradient descent, leading to more personalized models for each user.
@@
To achieve generalizability and adaptability, Multi-Task Representation Learning (MTRL) has emerged as a promising methodology in machine learning, aiming to address multiple related tasks/clients simultaneously by leveraging shared representations [7]. The core principle of MTRL is based on the idea that, even though data in federated settings is often non-i.i.d. (non-independent and identically distributed) across clients, there exists a shared learned representation across multiple clients that effectively captures the common structural elements inherent to each client. This shared foundation enables the use of simple linear predictors in the final model layer, which are then fine-tuned to address the unique variations and requirements of each client. This approach significantly enhances the efficiency of model training processes, as it allows for the transfer and utilization of learned features across different but related tasks. Several works have studied MTRL to address different aspects of learning and model performance. Yu Xie et al. solved the problem of simultaneously optimizing node classification and link prediction tasks by applying a shared feature extraction module followed by client-specific modules [8]. Yang et al. proposed a deep MTRL framework using tensor factorization, which enables automatic knowledge sharing across all layers of a deep neural network. This method applies to both homogeneous and heterogeneous settings and eliminates the need for manual specification of multi-task sharing strategies [9]. In addition, recent studies [3, 10, 11, 12] have focused on the theoretical understanding of MTRL, particularly when the shared model exhibits specific structures, such as a low-rank representation.
@@
To address the challenge of Byzantine attacks in federated learning, Gouissem et al. developed a collaborative cross-check mechanism that enhances the system’s resilience by allowing clients to collaboratively identify malicious nodes based on each other’s contributions to model updates [13]. While effective, this approach introduces significant computational overhead due to the extensive cross-validation among clients, which can lead to inefficiencies. In contrast, Feng Lin et al. tackled Byzantine resilience using robust aggregation techniques, such as geometric median, Krum, and h-Krum, to address the matrix completion problem while ensuring resilience against Byzantine failures [14]. Recently, Li et al. studied Byzantine resilient FL for non-iid data by obtaining an inversed artificial gradient [4]. Zhang et al. investigated communication-efficient approaches for Byzantine-resilient FL and proposed a projected stochastic gradient descent (SGD) algorithm to enhance system robustness [15]. Their approach employed a Huber function-based robust aggregation with an adaptive threshold selecting strategy at the server to reduce the effects of Byzantine attacks.
@@
Despite these advancements, a unified approach for Byzantine-resilient personalized FL with strong generalization remains unavailable. In this paper, we introduce the Byzantine-resilient Multi-Task Representation Learning (BR-MTRL) framework to jointly address three critical challenges in FL: data heterogeneity, generalization, and Byzantine resilience. By incorporating geometric median and Krum aggregation methods into personalized FL, we enhance the global model’s resilience against Byzantine attacks. Our framework ensures that the models learned across distributed agents remain robust, reliable, and secure, even in the presence of adversarial influences, improving both performance and safety.
@@ I-A Contributions
  • We propose a Byzantine-resilient MTRL algorithm based on personalized federated learning. In our approach, all clients share a common representation while maintaining their local personalized heads. We introduce an alternating gradient descent approach that alternates between solving for the shared model and fine-tuning each client’s personalized head. To ensure robustness, our method employs Geometric Median (GM) and Krum-based mechanisms for aggregating client updates at the server, allowing the clients to learn a Byzantine-resilient shared representation while preserving task-specific personalizations.
  • We implemented our proposed approach in a truly federated environment by deploying it on a simulation testbed built on the Amazon Web Services (AWS) platform. We performed simulations on datasets, including CIFAR-10 and FEMNIST, and validated the effectiveness of BR-MTRL in mitigating the impact of Byzantine attacks while maintaining a robust learning performance.
@@ I-B Related Work
Personalized Federated Learning (PFL) has received a lot of research interest in recent years. Liang et al. [16] proposed the Local Global Federated Averaging (LG-FedAvg) method to address data heterogeneity, which focuses on learning low-dimensional features from clients’ data. Their approach emphasizes learning multiple local representations and a single global head. On the other hand, Arivazhagan et al. [17] proposed FedPer algorithm that explicitly divides the model architecture into base layers, trained across all devices via federated averaging and personalization layers, which are tailored locally on each device using its specific data. This structure not only promotes personalization by adapting the model to local data characteristics but also maintains a cohesive model performance across diverse environments by leveraging a shared base learning across all clients. In addition, Collins et al. [3] proposed the FedRep algorithm focusing on learning a global model shared across clients while allowing each client to maintain a personalized head model, similar to [17]. The algorithms in [3, 17] allow clients to collaborate on learning shared representations while maintaining personalized models tailored to their specific data distribution. [3, 17] are closely related to our work; however, they focus on PFL without considering the challenges posed by Byzantine attacks.
@@
To counter Byzantine attacks, many PFL systems incorporate robust aggregation techniques at the server, such as the geometric median and its variations. For example, Li et al. introduced AutoGM, an auto-weighted geometric median aggregation rule designed to enhance robustness in the presence of Byzantine failures, specifically model and data poisoning attacks [18]. AutoGM improves the classic geometric median by automatically excluding extreme outliers and re-weighting the remaining parameter updates based on a skewness threshold, offering a more adaptive and flexible solution. Similarly, Wang et al. proposed WGM-dSAGA, a federated learning strategy that employs a weighted geometric median for robust aggregation against Byzantine attacks [19]. The strategy incorporates COPOD, a parameter-free outlier detection technique, to identify and assign weights to client updates based on their potential maliciousness. This method effectively mitigates the influence of Byzantine clients on the global model through the weighted geometric median. Meanwhile, He et al. tackled Byzantine attacks by introducing a novel projected stochastic block gradient descent method [20]. Unlike the GM-based approaches, their strategy employs a server aggregation technique that combines local models using a Huber function-based descent step to manage the impact of malicious updates from Byzantine clients. While these studies have made significant strides in mitigating Byzantine attacks, they primarily focus on enhancing aggregation techniques on the server side. In contrast, our work introduces a Byzantine-resilient MTRL framework that learns and leverages a shared representation to enhance the quality of each client’s model, while ensuring both transferability and resilience.
@@ II-A System Model
Our scheme comprises $n$ clients and a central server, with a subset of clients being malicious. The system model of the proposed MTRL scheme is presented in Figure 1. In each iteration round $t$ the proposed scheme follows these steps: (1) Each client fine-tunes and updates its local head $h^{t}_{i}$ using local data. (2) Clients update their estimates of the shared representation $\phi^{t}_{i}$. (3) Clients upload their updates to the server. (4) The server applies geometric median-based robust aggregation to compute the new shared representation $\phi^{t}$. (5) The updated representation is sent back to all clients.
@@
Malicious clients. We consider two types of clients: Byzantine and honest. Byzantine clients manipulate the model sent to the central server by providing malicious model updates, aiming to corrupt the final aggregation. In contrast, honest clients contribute genuine updates, ensuring reliable and accurate training. We consider a white-box attack scenario, where the adversary has the capability to poison the global model.
Robust server. To minimize the impact of Byzantine clients, we apply the Geometric Median (GM) and Krum to the weights received from clients, which is then transmitted to the server. These techniques offer a more resilient aggregation by effectively reducing the influence of outlier updates. By using the median instead of the mean, GM reduces outlier distortion, ensuring the global model reflects honest client updates. Krum enhances resilience by selecting the client update that is most similar to its nearest neighbors, thereby mitigating the impact of malicious updates on the global model.
@@ II-B Design Goals
Byzantine robustness. The proposed scheme learns an unbiased, personalized model, which minimizes the impact of malicious model updates.
Representation learning and accuracy. The proposed scheme must learn the shared model and the local client head and needs to ensure that the range of accuracy is acceptable.
Transferability. The proposed scheme should generalize to new, unseen clients by leveraging the learned shared model while training only the local client head, enabling efficient learning even with limited data.
@@ III Proposed Byzantine Resilient MTRL Algorithm
Client Update. In each round $t\in[T]$, we select a fraction $\alpha\in(0,1]$ of clients to participate in the client update. During this update, client $i$ performs ${\tau_{h}}$ local gradient steps to optimize its head model based on the current global representation ${\phi^{t-1}}$, which is shared by the server. Specifically, for each step $s=1,\dots,{\tau_{h}}$, client $i$ updates its head model as follows:
@@
Server Update. Once the local updates for the head and shared representation are completed, the client participates in the server update by sending its locally updated representation $\phi_{i}^{t,\tau_{\phi}}$ to the server. Prior to sending the updates, Byzantine clients manipulate the local representations by adding noise,
@@
where $\sigma$ is a scalar and $\mathcal{N}(0,I)$ denotes standard Gaussian noise. All clients $i\in[n]$ sends their local estimate of the representation, $\phi_{i}^{t,{\tau_{\phi}}}$, to the central server.
Robust Aggregation.To ensure robust aggregation, the server calculates Geometric Median for each layer of the shared model by finding a matrix $P$ by solving the following optimization
${\mathrm{GM}}(\phi_{1}^{t,\tau_{\phi}},\cdots,\phi_{n}^{t,\tau_{\phi}})=\arg\min_{P\in\mathbb{R}^{d\times k}}\sum_{i=1}^{n}\|\phi_{i}^{t,\tau_{\phi}}-P\|_{F},$
@@
For $k\neq k^{\prime}$, $k\rightarrow k^{\prime}$ is the set of $n-f-2$ nearest neighbors of $\phi_{k}^{t,\tau_{\phi}}$, where $f$ is the number of Byzantine clients. Thus $\phi_{k^{*}}^{t,\tau_{\phi}}$ is the model with minimal summed distance to its neighbors.
@@
Parameters: Participation rate $\alpha$, step size $\eta$, number of local updates for the head $\tau_{h}$ and for the global representation $\tau_{\phi}$, number of communication rounds $T$.
@@
2:     Server receives a batch of clients $I^{t}$ of size $rn$
3:     Server sends current representation $\phi^{t-1}$ to these clients
4:     for each client $i\in I^{t}$ do
5:         Client $i$ initializes $h_{i}^{t,0}$, $h_{i}^{t,0}\leftarrow h_{i}^{t-1,\tau_{h}}$
6:         for $s=1$ to $\tau_{h}$ $\triangleright$ Client $i$ makes $\tau_{h}$ head updates
7:              $h_{i}^{t,s}\leftarrow\text{GD}(f_{i}(h_{i}^{t,s-1},\phi^{t-1}),h_{i}^{t,s-1},\eta)$
8:         end for
9:         Client $i$ initializes $\phi_{i}^{t,0}$ $\phi_{i}^{t,0}\leftarrow\phi^{t-1}$
10:         for $s=1$ to $\tau_{\phi}$ $\triangleright$ Client $i$ makes $\tau_{\phi}$ updates to $\phi$
11:              $\phi_{i}^{t,s}\leftarrow\text{GD}(f_{i}(h_{i}^{t,\tau_{h}},\phi_{i}^{t,s-1}),\phi_{i}^{t,s-1},\eta)$
12:         end for
13:         if Client $i$ is malicious then
14:              $\phi_{i}^{t,\tau_{\phi}}=\phi_{i}^{t,\tau_{\phi}}+\sigma\cdot\mathcal{N}(0,I)$
15:         end if
16:         Client $i$ sends updated representation $\phi_{i}^{t,\tau_{\phi}}$ to server
17:     end for
18:     for each client $i\notin I^{t}$ do
19:         Set $h_{i}^{t,\tau_{h}}\leftarrow h_{i}^{t-1,\tau_{h}}$
20:     end for
@@
22:end for

### v3

@@ Abstract
In this paper, we propose BR-MTRL, a Byzantine-resilient multi-task representation learning framework that handles faulty or malicious agents. Our approach leverages representation learning through a shared neural network model, where all clients share fixed layers, except for a client-specific final layer. This structure captures shared features among clients while enabling individual adaptation, making it a promising approach for leveraging client data and computational power for personalized federated learning. To learn the model, we employ an alternating gradient descent strategy: each client optimizes its local model, updates its final layer, and sends estimates of the shared representation to a central server for aggregation. To defend against Byzantine agents, we employ two robust aggregation methods for client-server communication, Geometric Median and Krum. Our method enables personalized learning while maintaining resilience in distributed settings. We implemented the proposed algorithm in a federated testbed built using Amazon Web Services (AWS) platform and compared its performance with various benchmark algorithms and their variations. Through experiments using real-world datasets, including CIFAR-10 and FEMNIST, we demonstrated the effectiveness and robustness of our approach and its transferability to new unseen clients with limited data, even in the presence of Byzantine adversaries.
@@ I Introduction
Machine learning has traditionally relied on centralized settings, where a single model is trained on large datasets stored in one central location. This approach requires gathering data from all clients into a central data center, which raises significant data privacy concerns. Federated Learning (FL) addresses these challenges by using a distributed framework that preserves data privacy while overcoming the problem of isolated data sources [1]. FL enables multiple clients, such as mobile devices or institutions, to collaboratively train a global model without sharing their raw data. A central server is responsible for coordinating the clients by aggregating the models received from each client and broadcasting the updated model back to them in each round of communication. McMahan et al. introduced the Federated Averaging (FedAvg) algorithm, which facilitates the training of a single shared global model designed to achieve optimal performance on average across all participating clients [1]. The approach offers several advantages, such as simplicity in implementation and a streamlined aggregation process, making it scalable for large client populations. Additionally, the global model promotes uniformity, ensuring consistent performance for clients with similar data distributions. However, FL approaches encounter three primary bottlenecks: (i) Personalization: In real-world scenarios, collaborating client tasks are typically ‘different yet related’, making a single global model suboptimal for all tasks. Instead, personalized models are better suited to capture the unique characteristics and requirements of each client. (ii) Transferability: Federated models often struggle to generalize to unseen clients that were not part of the training process, limiting their effectiveness in dynamic, real-world applications. (iii) Byzantine attacks: FL is particularly vulnerable to Byzantine attacks, as it relies on distributed client updates; a faulty or malicious client can potentially corrupt the global model, compromising performance and reliability.
@@
Recent works proposed personalized FL algorithms. Deng et al. proposed the Adaptive Personalized Federated Learning (APFL) algorithm by combining local and global model optimization [2]. This method allows clients to train personalized models that better generalize on local distributions while still contributing to the learning of a robust global model. Fallah et al. proposed a method called Personalized Federated Learning via Model-Agnostic Meta-Learning (MAML) by applying a meta-learning framework [3]. This method allows clients to quickly adapt a shared initial model to their local data distributions by performing a few steps of gradient descent, leading to personalized models for each user.
@@
To achieve generalizability, Multi-Task Representation Learning (MTRL) has emerged as a promising methodology, aiming to address multiple related tasks/clients simultaneously by leveraging shared representations [4]. The core principle of MTRL is based on the idea that, even though data in federated settings is often non-i.i.d. (non-independent and identically distributed) across clients, there exists a shared learned representation across multiple clients that effectively captures the common structural elements inherent to each client. This shared foundation enables the use of simple linear predictors in the final model layer, which are then fine-tuned to address the unique variations and requirements of each client. This approach significantly enhances the efficiency of model training processes, as it allows for the transfer and utilization of learned features across different but related tasks. MTRL has demonstrated remarkable success across various applications such as natural language processing domains GPT-2, GPT-3, Bert, vision domains CLIP, and robotics. Recent studies [5, 6, 7, 8] have focused on the theoretical understanding of MTRL, particularly when the shared model exhibits specific structures, such as a low-rank representation.
@@
To address Byzantine attacks in federated learning, Gouissem et al. developed a collaborative cross-check mechanism by allowing clients to collaboratively identify malicious nodes based on each other’s contributions to model updates [9]. While effective, this approach introduces significant computational overhead due to the extensive cross-validation among clients. In contrast, Lin et al. tackled Byzantine resilience using robust aggregation techniques, such as geometric median, Krum, and h-Krum, to address the matrix completion problem [10]. Recently, Li et al. studied Byzantine resilient FL for non-iid data by obtaining an inversed artificial gradient [11]. Zhang et al. investigated communication-efficient approaches for Byzantine-resilient FL and proposed a projected stochastic gradient descent (SGD) algorithm to enhance system robustness [12]. Their approach employed a Huber function-based robust aggregation with an adaptive threshold selecting strategy at the server to reduce the effects of Byzantine attacks.
@@
Despite these advancements, a unified approach for Byzantine-resilient personalized FL with transferability remains underexplored. In this paper, we introduce the Byzantine-resilient Multi-Task Representation Learning (BR-MTRL) framework to jointly address three critical challenges in FL: personalization, transferability, and Byzantine resilience. By incorporating geometric median and Krum aggregation methods into the shared representation, while updating the task-specific parameter locally we enhance the global model’s resilience against Byzantine attacks. Our framework ensures that the models learned across distributed agents remain robust, reliable, and secure, even in the presence of adversarial influences, improving both performance and safety.
@@ I-A Contributions
  • We propose a BR-MTRL algorithm for personalized federated representation learning. In our approach, all clients share a common representation while maintaining their local personalized heads. We introduce an alternating gradient descent approach that alternates between solving the shared model and fine-tuning the task-specific head of each client locally. To ensure robustness, our method employs Geometric Median (GM) and Krum-based mechanisms to aggregate client updates on the server, allowing clients to learn a Byzantine-resilient shared representation while preserving task-specific personalizations.
  • We implemented our proposed approach in a truly federated environment by deploying it on a simulation testbed built on the Amazon Web Services (AWS) platform. We performed simulations on datasets, including CIFAR-10 and FEMNIST on the AWS testbed, and validated the effectiveness of BR-MTRL in mitigating the impact of Byzantine attacks while maintaining learning performance.
@@ I-B Related Work
Personalized Federated Learning (PFL) has received a lot of research interest in recent years. Liang et al. [13] proposed the Local Global Federated Averaging (LG-FedAvg) method, which focuses on learning low-dimensional features from clients’ data. Their approach emphasizes learning multiple local representations and a single global head. On the other hand, Arivazhagan et al. [14] proposed FedPer algorithm that explicitly divides the model architecture into base layers, trained across all devices via federated averaging and personalization layers, which are tailored locally on each device using its specific data. This structure not only promotes personalization by adapting the model to local data characteristics but also maintains a cohesive model performance across diverse environments by leveraging a shared base learning across all clients. In addition, Collins et al. [5] proposed the FedRep algorithm focusing on learning a global model shared across clients while allowing each client to maintain a personalized head model, similar to [14]. The algorithms in [5, 14] allow clients to collaborate on learning shared representations while maintaining personalized models tailored to their specific data distribution, similar to our approach; however, they focus on PFL without considering the challenges posed by Byzantine attacks.
@@
To counter Byzantine attacks, many PFL systems incorporate robust aggregation techniques at the server, such as the geometric median (GM) and its variations. For example, Li et al. introduced AutoGM, an auto-weighted GM aggregation rule designed to enhance robustness in the presence of Byzantine failures, specifically model and data poisoning attacks [15]. AutoGM improves the classic GM by automatically excluding extreme outliers and re-weighting the remaining parameter updates based on a skewness threshold, offering a more adaptive and flexible solution. Similarly, Wang et al. proposed WGM-dSAGA, a weighted GM for robust aggregation [16]. The strategy incorporates COPOD, a parameter-free outlier detection technique, to identify and assign weights to client updates based on their potential maliciousness. Meanwhile, He et al. tackled Byzantine attacks by introducing a novel projected stochastic block gradient descent method [17]. Unlike the GM-based approaches, their strategy employs a server aggregation technique that combines local models using a Huber function-based descent step to manage the impact of malicious updates from Byzantine clients. While these studies have made significant strides in mitigating Byzantine attacks, they primarily focus on enhancing aggregation techniques on the server side. In contrast, our work introduces a Byzantine-resilient MTRL framework that learns and leverages a shared representation to enhance the quality of each client’s model, while ensuring both transferability and resilience.
@@ II-A System Model
Our scheme comprises $n$ clients and a central server, with a subset of clients being malicious. The system model of the proposed MTRL scheme is presented in Figure 1. In each iteration round $t$ the proposed scheme follows these steps: (1) Each client fine-tunes and updates its local head $h^{t}_{i}$ using local data. (2) Clients update their estimates of the shared representation $\phi^{t}_{i}$. (3) Clients upload their updates to the server. (4) The server applies geometric median or Krum-based robust aggregation to compute the new shared representation $\phi^{t}$. (5) The updated representation is sent back to all clients.
@@
Malicious clients. We consider two types of clients: Byzantine and honest. Byzantine clients attack the global model update in two different ways: 1) corrupt the model sent to the central server by providing malicious model updates and 2) provide incorrect labels for their data. The attacker’s goal is to corrupt the final learned model. In contrast, honest clients contribute genuine updates, ensuring reliable and accurate training.
Robust server. To minimize the impact of Byzantine clients, we apply the Geometric Median (GM) and Krum to the weights received from clients, which is then transmitted to the server. The aggregation at the server is for the shared model among all the clients, which motivates the selection of GM and Krum aggregation techniques. By using the median instead of the mean, GM reduces outlier distortion, ensuring the global model reflects honest client updates. Krum enhances resilience by selecting the client update that is most similar to its nearest neighbors, thereby mitigating the impact of malicious updates.
@@ II-B Design Goals
Representation learning and model accuracy. The proposed method must jointly learn a shared representation model and personalized (local) client/task specific heads, while ensuring acceptable accuracy across the tasks.
Transferability. The proposed scheme should generalize to new, unseen clients/tasks by leveraging the learned shared model while training only the local client head, enabling efficient meta learning even with limited data.
Byzantine robustness. The proposed scheme learns an unbiased, personalized model, which minimizes the impact of malicious model updates and mislabeled data.
@@ III Proposed Byzantine Resilient MTRL Algorithm
Client Update. In each round $t\in[T]$, all clients participate in the client update. During this update, client $i$ performs $\tau_{h}$} local gradient steps to optimize its head model based on the current global representation $\phi^{t-1}$}, which is shared by the server. Specifically, for each step $s=1,\dots,{\tau_{h}}$, client $i$ updates its head model as follows:
@@
Server Update. Once the local updates for the head and shared representation are completed, the client updates the server by sending its locally updated representation denoted by $\phi_{i}^{t,\tau_{\phi}}$ to the server. We consider two types of Byzantine behaviors among malicious clients. In the Scaled Random (SR) attack, Byzantine clients inject scaled random noise into their local model updates to distort the global aggregation.
@@
where $\sigma$ is a scalar and $\mathcal{N}(0,I)$ denotes standard Gaussian noise. In the Mislabeling (ML) attack, Byzantine clients switch the labels of their local training data to cause misclassification and misleading gradients. All clients $i\in[n]$ sends their local estimate of the representation, $\phi_{i}^{t,{\tau_{\phi}}}$, to the central server.
Robust Aggregation. To ensure robust aggregation, the server calculates GM for each layer of the shared model by finding a matrix $P$ by solving the following optimization
$\vskip-5.69054pt{\mathrm{GM}}(\phi_{1}^{t,\tau_{\phi}},\cdots,\phi_{n}^{t,\tau_{\phi}})=\arg\min_{P\in\mathbb{R}^{d\times k}}\sum_{i=1}^{n}\|\phi_{i}^{t,\tau_{\phi}}-P\|_{F},$
@@
For $\neq k^{\prime}, kk$$\rightarrow k^{\prime}$ is the set of nf2n-f-2 nearest neighbors of $\phi_{k}^{t,\tau_{\phi}}$, where $f$ is the number of Byzantine clients. Thus $\phi_{k^{*}}^{t,\tau_{\phi}}$ is the model with minimal summed distance to its neighbors.
@@
Parameters: step size $\eta$, number of local updates for the head $\tau_{h}$ and for the global representation $\tau_{\phi}$, number of communication rounds $T$.
@@
2:  Server sends current representation $\phi^{t-1}$ to all clients
3:  for each client $i=1,2,\dots,n$ do
4:   Client $i$ initializes $h_{i}^{t,0}$, $h_{i}^{t,0}\leftarrow h_{i}^{t-1,\tau_{h}}$
@@
6:     $h_{i}^{t,s}\leftarrow\text{GD}(f_{i}(h_{i}^{t,s-1},\phi^{t-1}),h_{i}^{t,s-1},\eta)$
7:   end for
8:   Client $i$ initializes $\phi_{i}^{t,0}$ $\phi_{i}^{t,0}\leftarrow\phi^{t-1}$
[…]
10:     $\phi_{i}^{t,s}\leftarrow\text{GD}(f_{i}(h_{i}^{t,\tau_{h}},\phi_{i}^{t,s-1}),\phi_{i}^{t,s-1},\eta)$
11:   end for
12:   if Client $i$ is malicious then
13:     $\phi_{i}^{t,\tau_{\phi}}=Attack(\phi_{i}^{t,\tau_{\phi}})$
14:   end if
15:   Client $i$ sends updated representation $\phi_{i}^{t,\tau_{\phi}}$ to server
16:  end for
17:  Server computes the new representation as $\phi^{t}\leftarrow\mathrm{GM}(\phi_{1}^{t,\tau_{\phi}},\cdots,\phi_{n}^{t,\tau_{\phi}})\mbox{~or~}\mathrm{Krum}(\phi_{1}^{t,\tau_{\phi}},\cdots,\phi_{n}^{t,\tau_{\phi}})$
18:end for